OpenStudy (anonymous):

Solve: 5c − c + 10 = 34

OpenStudy (anonymous):

c=6

OpenStudy (anonymous):

ty wat about this one Solve: 2(x + 1) + 4 = 8

OpenStudy (anonymous):

1

OpenStudy (anonymous):

c = 6

OpenStudy (anonymous):

ohh ty how about this Solve: y over negative 6 + 5 = 9

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

y = -9

OpenStudy (anonymous):

thats not one of the choices :/

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

if of course the y over 6 is in parenthesis

OpenStudy (anonymous):

yes 24 ty do u know this one> Solve 3c + 4 = −20

OpenStudy (anonymous):

-8 :D

OpenStudy (anonymous):

ty! these 2? 1.Solve: −5(x − 2) = 15 2.Solve: 2x + 36 + 40 = 180

OpenStudy (anonymous):

can i atleast get a gold medal here ? :D

OpenStudy (anonymous):

yes hehe

OpenStudy (anonymous):

1.) x= -1 2.) x=52

OpenStudy (anonymous):

god ur good lol last onoe ;d What is the first step in solving the equation −4x − 16 = 20?

OpenStudy (anonymous):

x= -9

OpenStudy (anonymous):

noo what is the first step you do to solve it

OpenStudy (anonymous):

are there any parenthesis that you are leaving out?

OpenStudy (anonymous):

oh nvm got it lol ty

OpenStudy (anonymous):

plug in the -9 where x is okay so it would be -4 * -9 = 36 36 - 16 = 20 :)